nyard and Barley's circus is in town, and Lorena is going with her father. When Lorena and her father arrive at the performance, they can choose from six entrances to the ground floorof the main arena. There are two stairways and one elevator that go to the balcony. How many different paths can Lorena and her father take from outside the arena to the balcony?
Found 3 solutions by ikleyn, greenestamps, math_tutor2020:
Answer by ikleyn(52946)

.

    6*2*1 = 12.    ANSWER

Solved.

--------------------

I wrote it so shortly, because it is more than obviously / evidently.


///////////////////


After finding several different interpretations by other tutors,
my advise to the author of this post and the problem's composer
is to throw his composition to the closest garbage bin and do not show it to people around.



Answer by greenestamps(13219)



It's interesting that the other tutor would say the solution is obvious, when the statement of the problem allows many different interpretations -- of which the interpretation she uses is not the most likely.

With the interpretation used by the other tutor, there are 6 choices of the entrance to use to get to the main floor; then to get to the balcony there are a choice of 2 stairways, followed by an elevator. The number of paths using that interpretation is then 6*2*1 = 12.

It is far more likely that there are 6 choices for the entrance to use to get to the main floor, and then there are the choices of 2 stairways OR an elevator to get to the balcony. With that interpretation, the number of paths is 6*3 = 18.

And another possibility -- very unlikely but possible with the way the problem is worded -- is that there are 6 choices for the entrance to use to get to the main floor, and then to get to the balcony you have to go up 2 stairways and then use an elevator. With that interpretation, there is only 1 way to get from the main floor to the balcony, so the number of paths is only 6*1 = 6.

ANSWER: 18 (probably; but other interpretations give different answers)


Answer by math_tutor2020(3821)


Answer: 18

Reason:
(6 entrances)*(3 methods to get to the balcony) = 18 pathways